May 1, 2012 Readout

From GlueXWiki
Jump to: navigation, search

Teleconference Time: 11:00 a.m. EST

  • ESNET (Number is 8542553) and EVO session (GlueX Calorimetry meeting room)
  • Phone connection only upon request.
    • +1-866-740-1260 : US and Canada
    • +1-303-248-0285 : International
    • then use participant code: 3421244# (the # is needed when using the phone)
    • OR from type access code 3421244 into "join a meeting" without the # (you need java plugin)

Participant Direct Lines

  • JLab Phone: in CC F326 is 757-269-6337 (usual room)
  • JLab Phone in CC L207 is 757-269-7084
  • Phone in the Regina Videoconference Suite is 306-585-4204
  • Athens Phone: in Christina's office is 011-30-210-727-6947

Items for followup from previous meeting

  1. Beamtest
    • Draft a run plan (Elton and Zisis)
  2. SiPMs
    • Plot dark rate of SiPM measurements under the same conditions as for Hamamatsu for each of comparison (Orlando and Serguei)

Presentations from recent meetings

Tentative Agenda

  1. Announcements
  2. Review of Action Items
  3. Talks at the Collaboration Meeting May 21-23.
  4. BCAL Beam Tests 2012
    • Updates to readout board (Fernando)
  5. Update from USM
    1. USM Web Interface to DB
  6. Status of Testing at JLab
    1. Hall-D ELOG
  7. Other


Attending: Elton, Beni, John, Yi (JLab); Will (USM); Andrei (UofR)

  1. Announcements
  2. Review of Action Items
  3. Talks at the Collaboration Meeting May 21-23.
    • Talks at the collaboration meeting: Update on SiPMs and Light guides (possibly Javiera and Iñaki; or Will)
  4. BCAL Beam Tests 2012
    1. The electronics on the downstream end of the Bcal (wedge board) was replaced and has much reduced ringing.
    2. Data a week ago was 3 pass (2.541 GeV) and changed to 5 pass (5.551 GeV)
    3. Preliminary Analysis (Andrei). See Beam test meeting notes
      • Data from new board (downstream end) does not show nearly as much ringing as before (or the upstream side).
      • A week ago the spectra was very broad (3x expected width)
      • Data with one new board much cleaner (still wide, 2x expected width)
      • Mean position of peaks is about 700 (out of 6000) channels higher than expected from scaling energy changes. Problems could be due to analysis, pedestal subtraction (difficult due to ringing), or temperature difference between measurements (but should be compensated by circuit).
    4. Hall will be open tomorrow. We will replace chiller and change mini-Bcal configuration from normal incidence to 20 deg angle.
  5. Update from USM
    1. MPP Status
      • Production is now smooth
      • Almost ready to send tested batches. 2 or 3 MPP problems that need retesting.
      • Third shipment of MPPCs from Hamamatsu passed customs and is being delivered to the university.
    2. Light Guides
      • Project is moving nicely. Machined about 100 guides to date
      • Machining 5 surfaces with the 5-axis machine is straight-forward. The sixth surface is more complicated, but is going well. Jigs were built to hold the light guides on five (non-parallel) sides using friction and a pin in the LED hole. The machining is in production
      • The guides are made 17 microns oversized and reduced during the polishing process.
      • The polishing is not yet in production.
      • Preparing for 3 shifts per day to make up lost time. Still very labor intensive and need to pay attention to worker safety during the operation.
      • All parts in hand. Polishing setup is in the same room as the machining.
      • Aim to catch up and complete deliveries by the scheduled completion date.
    3. JLab promised to send the SiPM extractor and light guide gauges to Chile
      • The SiPM extractor is being hand-carried by Hayk
      • The templates (one for JLab and one for USM) are being produced and should be ready in about a week. Sergey may be able to take them back with him.
  6. Status of Testing at JLab
    1. Hall-D ELOG
      • Adjusted spacers between SiPM and plate to improve thermal contact
      • Technical help has been assigned (Bobby Bunton).
      • Expect to test about 50-80 samples per day. Presently testing at 15 deg C. Humidity in black box is kept low using si gel boxes. The dew point is 10 deg C
  7. Other